You have a portfolio that is invested 15 percent in Stock R, 34 percent in Stock S, and the remainder in Stock T. The beta of Stock R is .60, and the beta of Stock S is 1.15. The beta of your portfolio is 1.26. What is the beta of the Stock T?
Answer #1

Investment in T=100-(15+34)=51%

Portfolio beta=Respective beta*Respective weight

1.26=(0.6*0.15)+(1.15*0.34)+(Beta of stock T*0.51)

1.26=0.481+(Beta of stock T*0.51)

Beta of stock T=(1.26-0.481)/0.51

=1.53(Approx).